OpenStudy (anonymous):

find and label the vertex and the line of symmetry of f(x)-1/2x^2

OpenStudy (anonymous):

is it f(x)=(-1/2)x^2 ?

OpenStudy (anonymous):

the genaeral formula of this parabola would be x^2=-4a.y ----> x^2=-2y the vertex is (0,0) the line of symmetry is the x-axis , i.e y=0

OpenStudy (dumbcow):

well the coefficients are: a = -1/2 b=0 c = 0 for line of symmetry x = -b/2a = 0 y = (-1/2)*0 = 0 vertex: (0,0)

OpenStudy (anonymous):

http://www.youtube.com/watch?v=ttKRX1Zy7LE